OpenStudy (anonymous):

Get-A-Move-On moving company charges $40 for the day plus $0.80 per mile to rent a small moving truck. Truckin' It moving company charges $50 for the day plus $0.50 per mile. Complete the tables for the system of linear equations where m is the number of miles and c is the total cost.

OpenStudy (texaschic101):

Get-A-Move-On : 40 + 0.80m = C Truckin' It : 50 + 0.50m = C Now I am not sure what kind of table you need....1 for each company ?

OpenStudy (texaschic101):

For example, lets say you drive 50 miles... Get-A-Move-On would charge : 40 + 0.80(50) = C 40 + 4 = C 44 = C Truckin' It would charge : 50 + 0.50(50) = C 50 + 2.50 = C 52.50 = C so when driving 50 miles, Get-A-Move-On costs $44 and Truckin'It costs $52.50

OpenStudy (texaschic101):

are you understanding this ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

simple... first one....plug in 20 for m in the formulas I wrote and find what each company charges. GAMO would charge : 40 + 0.80(20) = C Truckin would charge : 50 + 0.50(20) = C the next one, sub in 40 for m and do the same thing... you getting how to do this ?

OpenStudy (texaschic101):

oh...wait...you only need results for Get-A-Move-On so sub in 20 for m 40 + 0.80(20) = C 40 + 1.60 = C 41.60 = C that goes in your first blank to find the others, just sub in the number for m into the formula and find C